Hi,
I can't seem to solve this problem...I have to find two consecutive, positive, even integers that the product of is 360. The original problem reads just like this: "The product of two consecutive positive even integers is 360. Find the integers."
Any help you can give me would be excellent! Thank you guys so, so, so much!!

Answer by stanbon(75887)

The product of two consecutive positive even integers is 360. Find the integers.
-------------
Let the even integers be (2x) and (2x+2)
(2x)(2x+2) = 360
4x(x+1) = 360
x(x+1) = 90
x^2+x-90=0
(x+10)(x-9) = 0
x=-10 or x=9
-----------
Positive means x=9
2x = 18 (1st even integer)
2x+2 = 20 (2nd even integer)
==============
Cheers,
Stan H.